# Graph Transformer Networks for Accurate Band Structure Prediction: An End-to-End Approach

### v1

@@ Discussion
In this discussion section, we will explore additional results, limitations of our model, and potential future directions for its development. Firstly, it is important to note that our model has limitations, including its inability to predict an unfixed and larger number of bands without prior user selection. To predict a variable number of bands, one approach could be to not fix $N_{b}$ and identify the maximum $N_{b}$ value, supplementing with empty virtual bands for materials with fewer actual bands. A similar method could be applied to the number of k-points. For the Materials Project dataset we used, the maximum values for $N_{b}$ and $N_{k}$ could be as high as 200 and 1500, respectively. Predicting a 200 by 1500 matrix would require substantial computational resources that exceed our current capabilities, but it is not unfeasible. Looking ahead, developing a large-scale model based on Transformer architecture to handle extensive band structure datasets is a promising research direction. Once such a model is trained, it could significantly reduce the computational resources needed for DFT calculations or Hamiltonian solving for a large datasets.
Figure 3: Predicted bands (red) vs real bands (black) on Materials Project band structure dataset. The dataset is divided into five equal sized subsets, with band energy prediction error low ($Q_{1}$) to high ($Q_{5}$).
@@
Another promising direction for future research involves the extraction of descriptors or targets for band structure-related properties that have more physical meaning, which could enable faster training and enhance model accuracy. In our current work, we applied FFT to convert raw band structure data from reciprocal space to an auxiliary ”frequency” space, treating the reciprocal space as ”time” space. This transformation produces targets with much smaller dimensions, which helps to reduce oscillations and overfitting in our model predictions. However, the exact physical meaning of this auxiliary space remains unknown, though it may be related to the localization of wave function in real space. Developing a low-dimensional descriptor for crystal band structures that can both decode from and encode into the band structure with minimal information loss would be extremely beneficial. Such a descriptor would simplify the representation of complex band structures, making it easier for models to learn and predict these structures accurately. This approach could significantly improve the performance of models designed for predicting band structures.
Methods
Graph transformer encoder
@@ Methods
For a given input crystal structure, we construct a crystal graph from the atoms in the primitive cell and their neighboring environment, selecting atoms within a predetermined searching radius. We consider 12 nearest neighbors within 8 Å searching radius of each atom to create the crystal graph for each input crystal structure. In the encoding phase, the crystal structure is initially represented as a graph $G$, where nodes represent atoms, and edges represents interatomic distances. Each vertex $i$ is characterized by a feature vector $\bm{x}_{i}^{0}$ based on properties of the atom such as group number, period number and atomic radius. The performance of model using this type of features is tested and compared to that using just one-hot encoding of atomic number as input features. Each edge $ij$ is represented by a feature vector $\bm{e}_{ij}$, which is a Gaussian function expansion of interatomic distances, which has been used frequently in previous works. We use a graph Transformer architecture within MPNN framework [31]. At each updating step $l$, edge feature $\bm{e}_{ij}$ and atom feature $\bm{x}_{i}^{l}$ are passed to linear layers without bias to generate edge embedding, query, key and value vectors: $\bm{m}_{ij}^{l}=W_{e}^{l}\bm{e}_{ij}$, $\bm{q}_{i}^{l}=W_{q}^{l}\bm{x}_{i}^{l}$, $\bm{k}_{i}^{l}=W_{k}^{l}\bm{x}_{i}^{l}$, $\bm{v}_{i}^{l}=W_{v}^{l}\bm{x}_{i}^{l}$, where $\bm{e}_{ij}$, $\bm{x}_{i}\in\mathbb{R}^{d}$ are $d$ dimensional feature vectors and $W\in\mathbb{R}^{d\times d}$ are $d\times d$ dimensional projection matrices.
@@
where $\langle\bm{q},\bm{k};\bm{m}\rangle=\frac{1}{\sqrt{d}}\bm{q}^{T}\bm{k}+\bm{m}$ is a modified multi-head scaled dot product with edge feature projection $\bm{m}$ included. The node features are updated by a skip connection first introduced in ResNet [23]: $\bm{x}_{i}^{l+1}=\bm{x}_{i}^{l}+\text{LayerNorm}\left(\sum_{j\in\mathcal{N}_{r}(i)}\alpha_{ij}^{l}\bm{x}_{j}^{l}\right)$. After $L$ iteration steps, a crystal feature is obtained by the global pooling of graph nodes $x^{L}=\sum_{i\in G}x_{i}^{L}$. This feature serves as the input for the graph2seq decoder.
Reciprocal space positional encoding
In decoding phase, we utilize a graph2seq model similar to Xtal2DoS, which has been used to predict electronic DOS. [8] The output from the encoder is utilized to generate key and value vectors. Each k-point along the k-path has coordinates $(k_{ix},k{iy},k_{iz})$. We featurize the k-point coordinates by using the positional encoding defined as:
$f_{n}(k_{i})=\exp{j2\pi\omega_{n}k_{i}}\$ (2)
@@
$\text{feat}=[\cos{(2\pi\omega_{0}k_{x})},...,\cos{(2\pi\omega_{\text{dim}-1}k_{x})},\sin{(2\pi\omega_{0}k_{x})},...,\sin{(2\pi\omega_{\text{dim}-1}k_{x})},...,\text{(repeat for {\hbox{y}} and {\hbox{z}} directions)}].$) ] . (4)
The similar form was been applied to real space coordinates in previous work and proven to be helpful in featurizing coordinate like input [26]. This initial feature is used for generating the query vector for each k-point along the k-paths. $x_{i}^{L}$ is used for creating key and value vectors. The initial k-point features, alongside the output from the graph Transformer decoder, are iteratively fed into the decoder for multiple time steps. Finally, the final output from the last layer is processed through a multi-layer perceptron (MLP).
Data availability
The processed band structure data for model training used in this study is available upon request.
Code availability
The code of the model used in this study is available upon request.
@@ Acknowledgements.
Competing interests
The authors declare no competing interests.

### v2

@@ II Discussion
In this section, we will explore additional results, limitations of our model, and potential future directions for its development. Firstly, it is important to note that our model has limitations, including its inability to predict an unfixed and larger number of bands without prior user selection. To predict a variable number of bands, one approach could be to not fix $N_{b}$ and identify the maximum $N_{b}$ value, supplementing with empty virtual bands for materials with fewer actual bands. A similar method could be applied to the number of k-points. For the Materials Project dataset we used, the maximum values for $N_{b}$ and $N_{k}$ could be as high as 200 and 1500, respectively. Predicting a 200 by 1500 matrix would require substantial computational resources that exceed our current capabilities, but it is not unfeasible.
@@
Another promising direction for future research involves the extraction of descriptors or targets for band structure-related properties that have more physical meaning, which could enable faster training and enhance model accuracy. In our current work, we applied FFT to convert raw band structure data from reciprocal space to an auxiliary ”frequency” space, treating the reciprocal space as ”time” space. This transformation produces targets with much smaller dimensions, which helps to reduce oscillations and overfitting in our model predictions. Looking ahead, developing a large-scale model based on transformer architecture to handle extensive band structure datasets is a promising research direction. Once such a model is trained, it can be directly deployed to make predictions and reduce the computational cost for DFT calculations or Hamiltonian solving for predicting band structures of unknown materials.
In conclusion, we have introduced Bandformer, the first end-to-end model that directly predicts electronic band structures from crystal structures. Our approach treats band structure prediction as a sequence translation task, using graph transformers to encode crystal structures and a graph2seq decoder to generate band energies along high-symmetry k-paths. By applying Fast Fourier Transform to band structures, we achieve efficient representation and stable training while maintaining physical accuracy. The model demonstrates strong performance on a diverse dataset from the Materials Project, accurately predicting both band energies and derived properties such as band gaps, band centers, and band dispersions. This work establishes that machine learning can effectively approximate the complex quantum mechanical mapping between crystal and electronic structures, enabling rapid electronic structure prediction for high-throughput materials screening and accelerating the discovery of functional materials with desired electronic properties.
III Methods
III.1 Graph transformer encoder
@@
For a given input crystal structure, we construct a crystal graph from the atoms in the primitive cell and their neighboring environment, selecting atoms within a predetermined searching radius. We consider 12 nearest neighbors within 8 Å searching radius of each atom to create the crystal graph for each input crystal structure. In the encoding phase, the crystal structure is initially represented as a graph $G$, where nodes represent atoms, and edges represent interatomic distances. Each vertex $i$ is characterized by a feature vector $\bm{x}_{i}^{0}$ based on properties of the atom such as group number, period number and atomic radius. The performance of model using this type of features is tested and compared to that using just one-hot encoding of atomic number as input features. Each edge $ij$ is represented by a feature vector $\bm{e}_{ij}$, which is a Gaussian function expansion of interatomic distances, which has been used frequently in previous works. We use a graph transformer architecture within MPNN framework [gilmer2017]. At each updating step $l$, edge feature $\bm{e}_{ij}$ and atom feature $\bm{x}_{i}^{l}$ are passed to linear layers without bias to generate edge embedding, query, key and value vectors: $\bm{m}_{ij}^{l}=W_{e}^{l}\bm{e}_{ij}$, $\bm{q}_{i}^{l}=W_{q}^{l}\bm{x}_{i}^{l}$, $\bm{k}_{i}^{l}=W_{k}^{l}\bm{x}_{i}^{l}$, $\bm{v}_{i}^{l}=W_{v}^{l}\bm{x}_{i}^{l}$, where $\bm{e}_{ij}$, $\bm{x}_{i}\in\mathbb{R}^{d}$ are $d$ dimensional feature vectors and $W\in\mathbb{R}^{d\times d}$ are $d\times d$ dimensional projection matrices.
@@
where $\langle\bm{q},\bm{k};\bm{m}\rangle=\frac{1}{\sqrt{d}}\bm{q}^{T}\bm{k}+\bm{m}$ is a modified multi-head scaled dot product with edge feature projection $\bm{m}$ included. The node features are updated by a skip connection first introduced in ResNet [22]: $\bm{x}_{i}^{l+1}=\bm{x}_{i}^{l}+\text{LayerNorm}\left(\sum_{j\in\mathcal{N}_{r}(i)}\alpha_{ij}^{l}\bm{x}_{j}^{l}\right)$. After $L$ iterations, a crystal feature is obtained by the global pooling of graph nodes $x^{L}=\sum_{i\in G}x_{i}^{L}$. This feature serves as the input for the graph2seq decoder.
III.2 Reciprocal space positional encoding
In decoding phase, we utilize a graph2seq model similar to Xtal2DoS, which has been used to predict electronic DOS. [8] The output from the encoder is utilized to generate key and value vectors. Each k-point along the k-path has coordinates $(k_{ix},k_{iy},k_{iz})$. We use the features for the k-point coordinates by passing k-point coordinates to the positional encoding defined as:
$f_{n}(k_{i})=\exp\left(j2\pi\omega_{n}k_{i}\right)\\$ (3)
@@ III.2 Reciprocal space positional encoding
. (5)
The similar form has been applied to real space coordinates in previous work and proven to be helpful in creating features for coordinates like input [25]. This initial feature is used for generating the query vector for each k-point along the k-paths. $x_{i}^{L}$ is used for creating key and value vectors. The initial k-point features, alongside the output from the graph transformer decoder, are iteratively fed into the decoder for multiple time steps. Finally, the final output from the last layer is processed through a multi-layer perceptron (MLP).
III.3 Real Fast Fourier Transform Representation
In this work, we apply a real Fast Fourier Transform (rFFT) to the positional-encoded sequence to compress the information for denser representation of band structures. After constructing the high-symmetry k-path and resampling to a fixed length of $N_{k}=128$, each k-point coordinate $(k_{x},k_{y},k_{z})$ is first mapped into a $d_{\text{model}}=256$ dimensional vector using the reciprocal space positional encoding described in the previous subsection. This results in an encoded representation of shape $(N_{k},d_{\text{model}})=(128,256)$.
We then apply the discrete rFFT implemented in PyTorch (torch.fft.rfft) along the $N_{k}$ dimension. The rFFT converts the input sequence from one dimensional subspace of reciprocal space into frequency space, producing an output of shape $(N_{k}/2+1,d_{\text{model}})=(65,256)$. Each column in this transformed representation encodes the amplitudes and phases of the oscillatory modes for a given embedding dimension.
Using rFFT provides several advantages. Firstly, the band structure values under LM scheme along continuous k-paths are smooth, leading to fast-decaying Fourier magnitudes. This property allows the model to concentrate on low-frequency modes that capture the essential trends of band dispersion. Second, the frequency-domain representation constrains the range of values, which stabilizes optimization and improves convergence during training. Third, rFFT effectively compresses the sequence representation from 128 to 65 points without significant information loss, reducing both computational cost and overfitting risk.
During model training, Bandformer predicts the real and imaginary components of the rFFT-transformed band structures. The predicted frequency-domain representation is then converted back to real space using the inverse rFFT, ensuring that the reconstructed band structures are directly comparable to those obtained from first-principles calculations.
IV Data availability
The processed band structure data for model training used in this study will be made available after publication of this work.
V Code availability
The code used in this study is available at https://github.com/qmatyanlab/Bandformer.
VI Competing interests
The authors declare no competing interests.